Alpha Company, on March 1, 2017 has a beginning Work in Process inventory of zero. All materials are added into production at the beginning of its production. There is only one production WIP inventory. On March 1, Alpha started into production 10,000 units. At the end of the month there were 5,000 units completed and transferred into the Finished Goods Inventory. The ending WIP was 60% complete with respect to conversion. For the month of March the following costs were incurred and recorded in the WIP:

            Direct Material                      $20,000

            Direct Labor                              8,000

            Factory Overhead                   32,000

Alpha uses the weighted-average process costing method. Use this information to determine for the month of March: (Round & enter final answers to the nearest penny.)

            1. the cost per equivalent unit of direct material

            2. the cost per equivalent unit of conversion

            3. the cost per unit transferred to finished goods
Answer #1

Calculation of Equivalents units of Direct materials and Conversion

Direct Material Conversion
Units started and completed 5000 5000
Ending work-in-process 5000 (5000*60%)= 3000
Equivalent units of Production 10000 8000

Units started and completed= Units started-Ending work-in-process units

= 10000-5000

= 5000 units

1) Cost per equivalent unit of direct material= Cost of Direct material/Equivalent units of Direct material

= $20000/10000= $2 per unit

2) Cost per equivalent unit of conversion= Cost of Conversion/Equivalent units of Conversion

= $(8000+32000)/8000= $5 per unit

3) Cost per unit transferred to finished goods= Cost per equivalent unit of direct material+Cost per equivalent unit of conversion

= $5+2= $7